# On the role of overparametrization in Quantum Approximate Optimization

### v1

@@ IV Conclusion
For both problem classes, we found that overparameterization serves as a sufficient condition for obtaining exact solutions with a high probability. For MAX-CUT on 2-regular graphs we further established that it also serves as a necessary condition — the overparameterization depth exactly matches the analytically determined optimal circuit depth, which we analytically prove to scale linearly with the problem size $n$ for even $n$. While this exact agreement does not hold for MAX-CUT instances on other graph classes ($k$-regular and random graphs), we observed an approximate agreement — most instances were solved either exactly at, or in close proximity to the overparameterization depth. The role of overparameterization gets further highlighted by the success probability, quantified as the fraction of optimization runs that converge to the optimum, which exhibited a pronounced increase upon transitioning from the underparametrized to the overparametrized regime. Taken together, these results indicate that, for MAX-CUT instances, the overparametrization depth offers a reasonable estimate of the optimal depth needed to achieve solutions with high probability.
@@
The MAX-2-SAT problem, despite having a Hamiltonian formulation reminiscent of MAX-CUT, exhibited markedly different behavior. A majority of instances were solved well within the underparametrized regime, $p/p_{c}\ll 1$. To put this in perspective, $\sim 85\%$ of the instances with $n=7$ were solved using circuits with depths as small as half the overparametrization depth. Likewise, most optimization runs converged deep in the underparametrized regime, resulting in a significantly high success probability even for $p/p_{c}\ll 1$. This behavior stands in stark contrast to the findings of [19], where the success probability remains negligible in the underparametrized regime and experiences an abrupt increase only when the number of circuit parameters approaches the dimension of the dynamic Lie algebra (DLA). It also differs markedly from related results for the Quantum Walk Optimization Algorithm (QWOA), a generalized version of QAOA, where it was shown that for all problems belonging to the class NP Optimization with Polynomially Bounded objective value (NPO-PB) but not belonging to the class BPPO (Bounded-error Probabilistic Polynomial Time Optimization) require an overparametrized QWOA sequence to be solved [46]. We believe these findings are largely encouraging. The fact that overparametrization is not necessary to solve MAX-2-SAT instances, highlights the potential of QAOA in the underparametrized regime. Moreover, this indicates that certain problems can be solved with substantially fewer resources than what would otherwise be needed if overparametrization were necessary.
V Acknowledgements
@@
Appendix A Combinatorial optimization problems: MAX-CUT and MAX-2-SAT
A.1 k-SAT
Boolean satifyability, or SAT, is the problem of determining weather a boolean formula written in conjunctive normal form (CNF) is satisfiable. It is possible to map any SAT instance via Karp reduction into $k$-SAT, which are restricted to $k$ literals per clause. In order to approximate solutions to SAT we embed the instance into a Hamiltonian as
$H_{\text{SAT}}=\sum_{j}P(j),$ (9)
where $j$ indexes clauses of an instance, and $P(j)$ is the tensor product of projectors that penalizes bit string assignments that do not satisfy the $j$-th clause. In this work we focused on the $k=2$ case.
A.2 MAX-CUT
Given a graph $(V,E)$, MAX-CUT is a problem of identifying the maximal cut of the graph, i.e. finding a graph bipartition with maximal number of edges between two components. Any instance of MAX-CUT problem can be encoded as an $n=|V|$ qubit problem Hamiltonian
$H=\sum_{(i,j)\in E}Z_{i}Z_{j},$ (10)
[…]
Appendix B Abrupt energy drop for the ring of dissagrees
@@
Appendix C Proof of Theorem 1
@@
Interestingly, in both cases of $n=4m$ and $n=4m+2$ the circuit is composed of multiple exponents with angles $\pi/4$, with the two distinct angles sitting exactly in the middle of the circuit—whether in the same layer or in the neighboring ones. Interestingly, some similar patterns in optimal parameter were observed for a few of $4$-regular graphs, yet no relation to the ring of disagrees was found. These results can be conveniently established using certain graphical rules for track Pauli backpropagation, which we leave for future works.
Appendix D Optimization strategy and data post-processing
There may be different heuristic strategies for optimization in QAOA [30]. In our work, we use a strategy similar to the one implemented in [45]. Namely, we start the optimization with $p=1$ ansatz layer. Once the optimization terminates, we use the optimal parameters $(\beta_{1}^{*},\gamma_{1}^{*})$ to initialize the optimization with $p=2$ layers at the parameters $(\beta_{1}^{*},\beta_{2}^{(0)},\gamma_{1}^{*},\gamma_{2}^{(0)})$, where $\beta_{2}^{(0)}$ and $\gamma_{2}^{(0)}$ are sampled uniformly. All the parameters are then optimized simultaneously. This procedure is repeated until reaching a desired number of layers $p$ or attaining the target accuracy.
Although this strategy may save computational time, much like any other initialization strategy, it has certain drawbacks. As each new layer is initialized randomly, the optimization with $p+1$ layers does not necessarily converge to lower cost function values compared to the $p$ layer optimization. That is, the relation $E_{p+1}^{*}(H)\leq E_{p}^{*}(H)$ may not hold for this strategy, as the optimization may stuck in a higher local minimum. Therefore, we post-process the data by setting $E_{p+1}^{*}(H)=E_{p}^{*}(H)$ whenever we obtained a higher energy error at $p+1$ layers compared to $p$ layers. This is demonstrated in figure 5, where we show the minimum energy error obtained for a MAX-CUT instance across $20$ optimization runs with and without the post-processing described above. Note that this post-processing is also applied in figure 2 in the main text.

### v2

@@ IV Conclusion and discussion
For both problem classes, we found that overparametrization serves as a sufficient condition for obtaining exact solutions with a high probability. For MAX-CUT on 2-regular graphs we further established that it also serves as a necessary condition — the overparametrization depth exactly matches the analytically determined optimal circuit depth, which we prove to scale linearly with the problem size $n$ for even $n$. This agreement approximatelly holds for other graph classes ($k$-regular and random graphs): most instances were solved either exactly at, or in close proximity to the overparametrization depth. The role of overparametrization gets further highlighted when considering the success probability, quantified as the fraction of optimization runs that converge to the optimum. In a good agreement with [36] the success probability exhibited a pronounced increase upon transitioning from the underparametrized to the overparametrized regime. Taken together, these results indicate that, for MAX-CUT instances, the overparametrization depth offers a reasonable estimate for the optimal depth needed to achieve solutions with high probability.
@@
The MAX-2-SAT problem, despite having a Hamiltonian formulation reminiscent of MAX-CUT, exhibited markedly different behavior. A majority of instances were solved well within the underparametrized regime, $p/p_{c}\ll 1$. To put this in perspective, $\sim 85\%$ of the instances with $n=7$ were solved using circuits with depths as small as half the overparametrization depth. Likewise, most optimization runs converged deep in the underparametrized regime, resulting in a significantly high success probability even at $p/p_{c}\sim 1/\sqrt{2^{n}}$. This behavior stands in a stark contrast to the findings of [36], where the success probability remains negligible and optimization stagnates in the underparametrized regime; the probability of finding the solution experiences an abrupt increase only when the number of circuit parameters approaches the dimension of the dynamical Lie algebra (DLA), i.e. when transitioning to the overparametrized region.
This behavior also differs markedly from related results for the Quantum Walk Optimization Algorithm (QWOA), a generalized version of QAOA, where it was shown that for all problems belonging to the class NP Optimization with Polynomially Bounded objective value (NPO-PB) but not belonging to the class BPPO (Bounded-error Probabilistic Polynomial Time Optimization) require a highly overparametrized QWOA sequence to be solved [15]. One arrives at a similar observation for the case of the unstructured search problem [5, 28] with QAOA; the dimension of DLA (and hence the overparametrization depth $p_{c}$) is upper bounded by $O(n^{2})$ due to the ansatz symmetries, while the optimal circuit depth is lower bounded by $O(\sqrt{2^{n}})$, due to the optimality of Grover scaling [28]. Nevertheless, we believe the findings of the current work are largely encouraging. The fact that overparametrization is not necessary to solve MAX-2-SAT instances highlights the potential of QAOA in the underparametrized regime.
The potential of the underparametrized circuits becomes further pronounced when noting that due to its reduced depth they are less prone to noise, making them more favorable for NISQ devices. Moreover, this indicates that certain problems can be solved with substantially fewer resources than what would otherwise be needed if overparametrization was necessary. Further improvement in performance can potentially be expected upon considering other modified versions of QAOA such as in [59]. Such investigations are pivotal as it could potentially allow one to counter the trainability limitations induced by barren plateaus — the phenomenon where the cost function gradients get exponentially suppressed. It has been observed that the overparametrization depth typically scales with the dimension of Hilbert space, i.e. exponentially with the system size, $p_{c}\propto 2^{n}$, while the gradients typically vanish already at the polynomial depth $p=poly(n)$ [19]. Thus, for large $n$ it is natural to expect that gradients become suppressed even in the underparametrized regime with normalized depth $p/p_{c}\ll 1$.
Acknowledgements
@@
Data availability
The data that support the findings of this article are openly available at [33].
Appendix A Abrupt energy drop for the ring of disagrees
@@
Appendix B Proof of Theorem 1
@@
Interestingly, in both cases of $n=4m$ and $n=4m+2$ the circuit is composed of multiple exponents with angles $\pi/4$, with the two distinct angles sitting exactly in the middle of the circuit—whether in the same layer or in the neighboring ones. Interestingly, similar patterns in optimal parameter were observed for a few of $4$-regular graphs, yet no relation to the ring of disagrees was found. These results can be conveniently established using certain graphical rules for track Pauli backpropagation, which we leave for future works.
Appendix C Optimization strategy and data post-processing
There may be different heuristic strategies for optimization in QAOA [57, 52, 53]. Among the most prominent are the ones described in [69]. There, the authors propose to reparametrize the ansatz angles via the so-called Fourier encoding, or interpolate the parameters found for $p$ ansatz layers to $p+1$ layers. Similarly to the latter, Ref. [38] introduces a bilinear initialization strategy for the $(p+1)$-th layer, which utilizes the parameters for the layers p1p-1 and $p$. More broadly, the optimal parameters can be transferred even between different problem instances within a problem class (e.g., between different graphs in the MAX-CUT problem) [13, 65].
While many optimization strategies for QAOA exist, we intend to test the performance of QAOA in relation to the overparametrization depth $p_{c}$, regardless of any heuristics applied. Therefore, the most appropriate strategy for our purposes would be using no heuristic strategy whatsoever. That is, ideally, we should have optimized each instance of the QAOA ansatz for each $p$ from scratch with random initializations. However, this would consume a lot more computational time. Therefore, in our work, we use a strategy similar to the one implemented in [4]. Namely, we start the optimization with $p=1$ ansatz layer. Once the optimization terminates, we use the optimal parameters $(\beta_{1}^{*},\gamma_{1}^{*})$ to initialize the optimization with $p=2$ layers at the parameters $(\beta_{1}^{*},\beta_{2}^{(0)},\gamma_{1}^{*},\gamma_{2}^{(0)})$, where $\beta_{2}^{(0)}$ and $\gamma_{2}^{(0)}$ are sampled uniformly. All the parameters are then optimized simultaneously. This procedure is repeated until reaching a desired number of layers $p$ or attaining the target accuracy.
To this end, the simple strategy of re-using the previous angles found for p1p-1 layers to initialize the $p$-layered QAOA serves as a compromise between saving computational resources and not introducing too much enhancement via heuristics. Although this strategy may save computational time, much like any other initialization strategy, it has certain drawbacks. As each new layer is initialized randomly, the optimization with $p+1$ layers does not necessarily converge to lower cost function values compared to the $p$ layer optimization. That is, the relation $E_{p+1}^{*}(H)\leq E_{p}^{*}(H)$ may not hold for this strategy, as the optimization may stuck in a higher local minimum. Therefore, we post-process the data by setting $E_{p+1}^{*}(H)=E_{p}^{*}(H)$ whenever we obtained a higher energy error at $p+1$ layers compared to $p$ layers. This is demonstrated in figure 5, where we show the minimum energy error obtained for a MAX-CUT instance across $20$ optimization runs with and without the post-processing described above. Note that this post-processing is also applied in figure 2 in the main text. Importantly, usage of more sophisticated heuristic training strategies might provide better starting points and thus improve convergence. This would only reduce the reported optimal depth $p^{*}$, making the underparametrized regime even more favorable.